Instructions: For each problem, complete the following parts: • Balance the chemical equation. • Identify the ‘given’ information (boxes 1-3) and the ‘needed’ answer (box 4-6) by the box number in the flowchart above. • Answer the question. Show your work and circle your answer

1. ____Al (s) +____O2 (g)à ____Al2O3 (s)

48 g of O2 produce how many grams of Al2O3?

Given Information: Box _____

Needed Answer: Box ______

2. ____NaCN (aq) + ____H2SO4 (aq) à ____HCN (g) + ____Na2SO4 (aq)

35 mL of 0.190 M H2SO4 produce how many moles of HCN?

Given Information: Box _____

Needed Answer: Box ______

3. ____Zn (s) + ____H2SO4 (aq) à ____ZnSO4 (aq) + ____H2 (g)

58.0 mL of 0.250 M H2SO4 produce how many grams of H2?

Given Information: Box _____

Needed Answer: Box ______
Answer #1

1) Balanced chemical reaction,

4Al + 3O2 ----> 2Al2O3

Moles of O2 = mass/molar mass = 48/32 = 1.5 moles

From reaction,

3 moles of O2 produce 2 moles Al2O3

So, 1.5 moles O2 will produce = 2*1.5/3 = 1 mole of Al2O3

Mass of Al2O3 = moles * molar mass = 1 * 102 = 102 grams ...Answer

2) 2NaCN + H2SO4 ----> Na2SO4 + 2HCN

Moles of H2SO4 = molarity * volume in liters = 0.190 * 35/1000 = 0.00665

Moles of HCN produced = 2 * 0.00665 = 0.0133 moles ...Answer

3) Zn + H2SO4 ----> ZnSO4 + H2

Moles of H2SO4 = 0.250 * 58/1000 = 0.0145 moles

From reaction,

1 mole H2SO4 produces 1 mole H2,

So, 0.0145 mole will produce = 0.0145 moles of H2

Mass of H2 = moles * molar mass = 0.0145 * 2 = 0.029 grams ....Answer